Find the length of the curve 3 v=ln(1 +t), 0< < 2. 1+ Length
Answer #1

ya mitt 10 (4)= vul-uri uz van een je dit र let). Slo-t-e (1485 (Let2 (L+t)(1)-HD (1+t2 dt caro = lett Itu da de (tt)? ادله (Let u=ttl du=dt 2 J SHilth? (lett dt = j Jituz du o d we have J fg = fg -ffg u f = July gl=112 g=-14 Jato 22 Juz! 2 JULHUZ941 th(dati +3 +3) -|-e + ln El +) = [+ Ion 16 +3)]- [-]+ bulv2+1} + (-10540935 3+ •B v2 46459) ---1-41421356270-88 373587) =